Air flows through a converging-diverging nozzle/diffuser. Assuming isentropic flow, air as an ideal gas, and constant specific heats determine the state at several locations in the system. Solve using equations rather than with the tables.
Note: The specific heat ratio and gas constant for air are given as k=1.4 and R=0.287 kJ/kg-K respectively.

--Given Values--
Inlet Temperature: T1 (K) = 321
Inlet pressure: P1 (kPa) = 588
Inlet Velocity: V1 (m/s) = 97
Area at nozzle inlet: A1 (cm^2) = 6.36
Throat area: A (cm^2) = 3.22
Area at diffuser exit: A (cm^2) = 4.35

a) Determine the Mach number at the inlet.
        Your Answer = Correct! Exact Answer= 0.2701 +/- 9.4E-04
b) Determine the stagnation temperature (K) at the inlet.
        Your Answer = Correct! Exact Answer= 325.68 +/- 1.4E+00
c) Determine the stagnation pressure (kPa) at the inlet.
        Your Answer = Correct! Exact Answer= 618.58 +/- 2.4E+00
d) Determine the theoretical throat area (cm^2) for sonic flow (A*).
        Your Answer = Correct! Exact Answer= 2.842 +/- 1.2E-02
e) Determine the Mach number at the throat.
        Your Answer = Incorrect.
f) Determine the temperature (K) at the throat.
        Your Answer =
g) Determine the velocity (m/s) at the throat.
        Your Answer =
h) Determine the pressure (kPa) at the throat.
        Your Answer =
i) Determine the mass flow rate (kg/s) through the nozzle
        Your Answer =
j) Determine the Mach number at the exit.
        Your Answer =
k) Determine the temperature (K) at the exit.
        Your Answer =
l) Determine the velocity (m/s) at the exit.
        Your Answer =
Answer #1

Solution we have V ma 0.2100 77.4.X28+X 321 :. madh number at inlet = 0.2700 (b) То с т у 2. 2Cp (972 ax1005 321 + = 325.68 (5. we have K/R-11 Ро Р [1+ Cread mar] =[ + 012.(0:270172] 14% 1.91 618.58 P press wie Pe 170.0.67 ek pa: at rie the rout : :1(K) we hare To 1 + (k-1) ma? TEK 325:68 1+ (0-2) (0,42227% ti T= 31.7.997 R. Temperature et the exit = 317.997k (l ma soma с

P.S: If you are having any doubt, please comment here, I will surely reply you.Please provide your valuable feedback by a thumbsup. Your thumbsup is a result of our efforts. Thanks in advance!